# Location-Dependent Cryptosystem

### v1

@@ Abstract
We present a location-dependent cryptosystem in which the decryption key is not transmitted as human- or machine-readable data, but implicitly encoded in precise time-of-flight differences of ultra-wideband (UWB) data transmission packets. The system leverages Ciholas DWETH101 hardware and a custom JMTK protocol to map a 32-byte SHA-256–derived AES key onto scheduled transmission timestamps. Only receivers located within a predefined spatial region can observe the packet timings that align with the intended “time slot” pattern, enabling them to reconstruct the key and decrypt the secret. Receivers outside the authorized region observe yield incorrect keys.
We implement a complete prototype that encrypts and transmits audio data using our cryptosystem, and only when the receiver is within the authorized data they are able to decrypt the data. Our evaluation demonstrates that the system (i) removes the need to share decryption passwords electronically or physically, (ii) ensures the decryption key cannot be recovered by the eavesdropper, and (iii) provides a non-trivial spatial tolerance for legitimate users.
@@ 1 Introduction
Cryptographic primitives such as AES, TDES, ECC, and ElGamal have made it possible to protect data from digital piracy and intellectual property theft [terra2016digitalpiracy, florunda2023infotheft]. These schemes are widely deployed across entertainment, cloud, financial, and research-driven industries [nist_sp80038a, ansix9_52_2016, pcidss_v4, holz2011surveyssl, koblitz1987ecccipher, elgamal1985publickey, menezes1996handbook]. However, while modern ciphers offer strong guarantees on confidentiality and integrity, they are fundamentally indifferent to the geographic location in which decryption occurs. Once an adversary obtains a valid decryption key, by intercepting it during transit or stealing it from an endpoint, there is nothing in the cryptosystem itself that prevents them from decrypting the protected content anywhere and at any time. These industries need new security measures to protect their intellectual property from corporate or international espionage.
@@
Therefore, a viable solution is that the user should be able to decrypt a file only if he or she is at an acceptable geographic location pre-authorized by the sender. Therefore, our system will transfer the encrypted key so that only the user at the authorized location can decrypt it, without having prior knowledge of the sender or the password. Therefore, the need for sharing the password electronically or physically is omitted.
Figure 1: Transmitters are used to generate the authorized region where the receiver can decrypt the file.
This intuition motivated our defense: designing a cryptosystem that ties decryption capability to a geographic region. Intuitively, we want a mechanism such that: (1.) a legitimate receiver located within an authorized region can reconstruct the decryption key automatically, without ever seeing the key in cleartext or manually entering a password, and (2.) any party outside that region, even if they capture all encrypted traffic, cannot recover the key from the signals they observe.
@@
We design and implement a location-dependent cryptosystem built on Ciholas DWETH101 [cuwb_anchors_2020] hardware and a custom JMTK protocol. The sender encrypts data with AES-256 [nist_fips197_aes] using a password provided by the sender, derives a fixed-length key (32 bytes) by hashing the password using SHA-256 [nist_fips1804_sha], and encodes that key across a sequence of 33 UWB transmissions. The mapping from password encrypted bytes to transmission timestamps is computed using:
@@
We implement a full prototype, where a UWB-based system encrypts audio files, transmits the key implicitly via packet timings, and automatically decrypts and plays the audio only within the authorized area. Our prototype demonstrates that:
@@ 2 Background
In industry, the encryption standard used to encrypt data is AES [nist_fips197_aes]. The encrypted data is then transmitted using any one of the cryptographic network protocols, such as Internet Protocol Security (IPsec) [rfc4301_ipsec]. However, the key of the encryption is independent of the location and it needs to be transferred to the receiver physically or electronically. Therefore, in our cryptosystem, to make sure the encrypted data can only be decrypted at the authorized location, the key of the encrypted data is associated with the approved location. The system will start the decryption process automatically without the need for the receiver’s assistance, thus protecting the password’s integrity.
@@ 3.1 Hardware
Our cryptosystem requires a device that can emit ultra-precise timing information. Therefore, we utilized Ciholas’ DWETH101 [cuwb_anchors_2020] (as shown in 2 as transmitter and NetApp server to relay DWETH101 timing information accurately up to eight nanoseconds. DWETH101 contains a DecaWave [qorvo_dw1000] chip that can emit timestamps accurately up to ten picoseconds. NetApp server uses the timestamp from DecaWave chip and makes it more precise by using the most significant bits to get information in the nanoseconds range or resulting in $\pm 3$ cm precision in space.
Figure 2: Transmitter modeules: DWETH101.
@@ 3.2.1 Audio Extraction
The software  [^1] The code is provided in https://github.com/kunmukh/Location_Dependent_Cryptosystem was developed in multiple part and as an example demonstration we chose audio signals. Therefore, the development of an audio extraction tool was important. The audio extraction script takes an audio file, input.wav, and converts it into principal frequency values. The values are then stored in a file, which can later be used to convert the principal frequency values back into audio. The script uses the ffmpeg utility to convert the .wav file to principal frequency values.
@@ 3.2.2 AES-256 and SHA-256 Implementation
We use AES-256 encryption, and the encryption process operates on a specific buffer (e.g., 32 bytes). AES-256 encrypts the buffer using an initialization vector (IV) and the provided key. AES encryption produces 32 bytes, or one block, of encrypted data from the unencrypted buffer. The encrypted block is stored in a text file that would be transmitted later. This encryption process continues until all data has been converted into encrypted blocks.
The decryption process takes the encrypted data and the password and recreates the unencrypted data. Our cryptosystem converts the user-provided password to a fixed 32-byte value using SHA-256. An SHA-256 hashing function has two important properties. First, it protects the integrity of the user’s password since no one has knowledge of the plain text key except the user. Secondly, it creates a fixed length key every time, thus broadening the collision domain of the encrypted passwords.
3.2.3 JMTK Protocol
The JMTK protocol uses the 32 bytes of the SHA password and transmits it via the timestamp difference. To transmit 32 bytes of SHA, 33 packets are needed, since the first transmission serves as a reference for the first SHA calculation.
@@ 3.2.3 JMTK Protocol
The JMTK protocol had to be made compatible with the NetApp. If the protocol asked the NetApp to transmit one of its packets when the NetApp was scheduled to transmit one of NetApp’s packets, then the JMTK packet would not be transmitted due to NetApp’s internal packet taking higher priority. Thus, an initial offset had to be introduced for the protocol to work. This ensured the protocol did not interfere with the scheduling of the NetApp’s internal packets and also marked the beginning of the 33-timestamp transmission sequence. The initial offset is called $T_{\text{startOffset}}$ and is numerically equal to 5 milliseconds or 319,488,000 NT ticks.
@@
Figure 3: Time window diagram.
@@
Figure 4: Slot window and Slot number calculation.
After determining all the constants, the JMTK protocol determines which values of the SHA need to be transferred. 5D will be used as an example. The transmission side will take the last transmission timestamp calculated, $T_{\text{tx}}(n-1)$, and add the time of flight of the packet from the last chosen anchor to the reception region, $T_{\text{distA}}(n-1)$. This gives the transmitter the reference point. Then, the transmission side adds $T_{\text{betweenOffset}}$. This ensures that the receiver is in the listening state. Then, the transmission side takes the SHA value SlotNumber 5D or 93 and multiplies it by $T_{\text{slot}}$. The final step is to subtract the time of flight, $T_{\text{distA}}(n)$, to account for the time it takes from this current packet to go from the current transmitter to the approved location. Therefore, the transmission equation is:
@@
Figure 5: Transmission and reception equation values on a timeline.
Figure 6: The eavesdropper and the cryptosystem.
@@
Figure 7: Example calculation showing the difference between the reference value calculated by the intended receiver and eavesdropper.
Figure 8: Example calculation showing the difference between the SHA calculated by the intended receiver and eavesdropper.
@@ 3.2.5 Cryptosystem Server and Client
3.3 Social, Political and Environmental Impact
There are no social or environmental concerns associated with this project. This is a research endeavor to create a secure and robust cryptosystem. This will benefit everyone and improve the social aspect of sharing. This project will make sharing secure and trustworthy. Ciholas hardware is FCC approved. As a result, it poses no environmental or health risks, as the electromagnetic radiation is within the approved limit.
There is a political concern with this project. If this project becomes completely robust with no vulnerabilities, various research and government agencies would like to gain access to it and try to prevent this information from spreading to other parties. Therefore, I will keep an online, up-to-date version of my project available  [^2] https://github.com/kunmukh/Location_Dependent_Cryptosystem, accessible from anywhere.
@@ 6 Conclusion
By carefully scheduled UWB packet timing, our location-dependent cryptosystem enforces that decryption is possible only where the sender intends. Legitimate receivers within the authorized zone reconstruct the correct key and transparently decrypt audio content, while receivers outside this zone, even if they capture all packets, observe shifted timing patterns that yield incorrect keys. The system operates without ever exposing the password in human- or machine-readable form on the wire and initiates decryption without explicit user action, reducing both operational friction and key-handling risk.

### v2

@@ Abstract
This paper presents a location-dependent cryptosystem in which the decryption key is not transmitted as human- or machine-readable data, but implicitly encoded in precise time-of-flight differences of ultra-wideband (UWB) data transmission packets. The system leverages Ciholas DWETH101 hardware and a custom TiCK (Timing-encoded Cryptographic Keying) protocol to map a 32-byte SHA-256–derived AES key onto scheduled transmission timestamps. Only receivers located within a predefined spatial region can observe the packet timings that align with the intended “time slot” pattern, enabling them to reconstruct the key and decrypt the secret. Receivers outside the authorized region observe yield incorrect keys.
A complete prototype is implemented that encrypts and transmits audio data using our cryptosystem, and only when the receiver is within the authorized data they are able to decrypt the data. Our evaluation demonstrates that the system (i) removes the need to share decryption passwords electronically or physically, (ii) ensures the decryption key cannot be recovered by the eavesdropper, and (iii) provides a non-trivial spatial tolerance for legitimate users.
@@ 1 Introduction
Cryptographic primitives such as AES, TDES, ECC, and ElGamal have made it possible to protect data from digital piracy and intellectual property theft [terra2016digitalpiracy, florunda2023infotheft]. These schemes are widely deployed across entertainment, cloud, financial, and research-driven industries [nist_sp80038a, ansix9_52_2016, pcidss_v4, holz2011surveyssl, koblitz1987ecccipher, elgamal1985publickey, menezes1996handbook, Mukherjee2023ProvNinja, Mukherjee2024ProvIoT, Mukherjee2025ProvDP, Mukherjee2025ProvSEEK, Mukherjee2025ZREx, Mukherjee2025ProvExplainer, wang2025provcreator]. However, while modern ciphers offer strong guarantees on confidentiality and integrity, they are fundamentally indifferent to the geographic location in which decryption occurs. Once an adversary obtains a valid decryption key, by intercepting it during transit or stealing it from an endpoint, there is nothing in the cryptosystem itself that prevents them from decrypting the protected content anywhere and at any time. These industries need new security measures to protect their intellectual property from corporate or international espionage.
@@
Therefore, a viable solution is that the user should be able to decrypt a file only if he or she is at an acceptable geographic location pre-authorized by the sender. Therefore, this system will transfer the encrypted key so that only the user at the authorized location can decrypt it, without having prior knowledge of the sender or the password. Therefore, the need for sharing the password electronically or physically is omitted.
Figure 1: Authorized-decryption geometry. Multiple UWB transmitters (anchors) define an authorized spatial region in which a receiver observes the intended time-of-flight–aligned packet timing pattern and can reconstruct the key for decryption.
This intuition motivated the defense presented in this paper: designing a cryptosystem that ties decryption capability to a geographic region. Intuitively, we want a mechanism such that: (1.) a legitimate receiver located within an authorized region can reconstruct the decryption key automatically, without ever seeing the key in cleartext or manually entering a password, and (2.) any party outside that region, even if they capture all encrypted traffic, cannot recover the key from the signals they observe.
@@
We design and implement a location-dependent cryptosystem built on Ciholas DWETH101 [cuwb_anchors_2020] hardware and a custom TiCK (Timing-encoded Cryptographic Keying) protocol. The sender encrypts data with AES-256 [nist_fips197_aes] using a password provided by the sender, derives a fixed-length key (32 bytes) by hashing the password using SHA-256 [nist_fips1804_sha], and encodes that key across a sequence of 33 UWB transmissions. The mapping from password encrypted bytes to transmission timestamps is computed using:
@@
We implement a full prototype, where a UWB-based system encrypts audio files, transmits the key implicitly via packet timings, and automatically decrypts and plays the audio only within the authorized area. The prototype demonstrates that:
@@ 2 Background
In industry, the encryption standard used to encrypt data is AES [nist_fips197_aes]. The encrypted data is then transmitted using any one of the cryptographic network protocols, such as Internet Protocol Security (IPsec) [rfc4301_ipsec]. However, the key of the encryption is independent of the location and it needs to be transferred to the receiver physically or electronically. Therefore, in this cryptosystem, to make sure the encrypted data can only be decrypted at the authorized location, the key of the encrypted data is associated with the approved location. The system will start the decryption process automatically without the need for the receiver’s assistance, thus protecting the password’s integrity.
@@ 3.1 Hardware
The cryptosystem requires a device that can emit ultra-precise timing information. Therefore, Ciholas’ DWETH101 [cuwb_anchors_2020] was used as transmitter and NetApp server to relay DWETH101 timing information accurately up to eight nanoseconds. DWETH101 contains a DecaWave [qorvo_dw1000] chip that can emit timestamps accurately up to ten picoseconds. NetApp server uses the timestamp from DecaWave chip and makes it more precise by using the most significant bits to get information in the nanoseconds range or resulting in $\pm 3$ cm precision in space.
@@ 3.2.1 Audio Extraction
The software  [^1] The code is provided in https://github.com/kunmukh/Location_Dependent_Cryptosystem was developed in multiple part and as an example demonstration an audio file was chosen, with the goal that the audio file will be decrypted only at the authorized located originally determined by the sender. Therefore, the development of an audio extraction tool was important. The audio extraction script takes an audio file, input.wav, and converts it into principal frequency values. The values are then stored in a file, which can later be used to convert the principal frequency values back into audio. The script uses the ffmpeg utility to convert the .wav file to principal frequency values.
@@ 3.2.2 AES-256 and SHA-256 Implementation
AES-256 encryption was used, and the encryption process operates on a specific buffer (e.g., 32 bytes). AES-256 encrypts the buffer using an initialization vector (IV) and the provided key. AES encryption produces 32 bytes, or one block, of encrypted data from the unencrypted buffer. The encrypted block is stored in a text file that would be transmitted later. This encryption process continues until all data has been converted into encrypted blocks.
The decryption process takes the encrypted data and the password and recreates the unencrypted data. The cryptosystem converts the user-provided password to a fixed 32-byte value using SHA-256. An SHA-256 hashing function has two important properties. First, it protects the integrity of the user’s password since no one has knowledge of the plain text key except the user. Secondly, it creates a fixed length key every time, thus broadening the collision domain of the encrypted passwords.
3.2.3 TiCK (Timing-encoded Cryptographic Keying) Protocol
The TiCK (Timing-encoded Cryptographic Keying) protocol uses the 32 bytes of the SHA password and transmits it via the timestamp difference. To transmit 32 bytes of SHA, 33 packets are needed, since the first transmission serves as a reference for the first SHA calculation.
@@ 3.2.3 TiCK (Timing-encoded Cryptographic Keying) Protocol
The TiCK protocol had to be made compatible with the NetApp. If the protocol asked the NetApp to transmit one of its packets when the NetApp was scheduled to transmit one of NetApp’s packets, then the TiCK packet would not be transmitted due to NetApp’s internal packet taking higher priority. Thus, an initial offset had to be introduced for the protocol to work. This ensured the protocol did not interfere with the scheduling of the NetApp’s internal packets and also marked the beginning of the 33-timestamp transmission sequence. The initial offset is called $T_{\text{startOffset}}$ and is numerically equal to 5 milliseconds or 319,488,000 NT ticks.
@@
Figure 2: NetApp timing window and network cadence. Illustration of the synchronization window in which the 33-packet TiCK transmission sequence is scheduled (starting at $T_{net}$$T_{startOffset}$
@@
Figure 3: Slotting scheme for timing-encoded key bytes. A SHA-256 byte is mapped to a discrete slot number (0–255) and transmitted via an inter-packet delay of SlotNumberSlotNumber x $T_{slot}$, with $T_{slot}$ chosen to match the desired spatial tolerance of the authorized region.
After determining all the constants, the TiCK protocol determines which values of the SHA need to be transferred. 5D will be used as an example. The transmission side will take the last transmission timestamp calculated, $T_{\text{tx}}(n-1)$, and add the time of flight of the packet from the last chosen anchor to the reception region, $T_{\text{distA}}(n-1)$. This gives the transmitter the reference point. Then, the transmission side adds $T_{\text{betweenOffset}}$. This ensures that the receiver is in the listening state. Then, the transmission side takes the SHA value SlotNumber 5D or 93 and multiplies it by $T_{\text{slot}}$. The final step is to subtract the time of flight, $T_{\text{distA}}(n)$, to account for the time it takes from this current packet to go from the current transmitter to the approved location. Therefore, the transmission equation is:
@@
Figure 4: Transmit/receive timeline for decoding one key byte. Timeline view of the transmission and reception equations showing how the receiver cancels fixed offsets (e.g., $T_{betweenOffset}$) and uses the remaining delay to recover the slot value (and thus the SHA byte).
Figure 5: Threat model: intended receiver vs. eavesdropper. An unauthorized receiver outside the approved region observes shifted time-of-flight timing, causing incorrect slot decoding and failure to reconstruct the correct key.
@@
Figure 6: Numerical example: reference mismatch outside the authorized region. Example calculation comparing the intended receiver’s reference timing to an eavesdropper’s; even small displacement changes the computed reference value used for slot decoding.
Figure 7: Numerical example: SHA-byte decoding divergence. Example showing that the intended receiver and eavesdropper decode different SHA values from observed timings, leading to different reconstructed AES keys and unsuccessful decryption outside the authorized region.
@@ 3.2.5 Cryptosystem Server and Client
3.3 Social and Environmental Impact
There are no social or environmental concerns associated with this project. This is a research endeavor to create a secure and robust cryptosystem. This will benefit everyone and improve the social aspect of sharing. This project will make sharing secure and trustworthy. Ciholas hardware is FCC approved, so the devices poses no environmental or health risks.
@@ 4 Results
5 Discussion
This study demonstrates that physical location can be elevated from an external policy constraint to a first-class input to key reconstruction by encoding a SHA-256, derived AES key into UWB packet timing patterns that are only decodable within an authorized region. While the prototype validates feasibility, several deployment and security questions naturally arise.
5.1 Design Trade-offs and Practical Constraints
A central design knob is the slot duration $T_{\text{slot}}$, which directly governs the spatial tolerance of the authorized region (e.g., a larger region implies a larger timing tolerance). In practice, this creates a location-utility trade-off: increasing tolerance reduces false rejections for legitimate receivers but also reduces the timing separation that protects against off-region reconstruction. The protocol also imposes real-time scheduling constraints (e.g., minimum inter-packet spacing so receivers can process packets), which affects throughput and robustness under load.
Does the security argument extend beyond a passive off-region eavesdropper?
The current rationale focuses on a passive adversary who records all packets but experiences different time-of-flight and therefore reconstructs an incorrect key. For stronger real-world assurance, the evaluation should explicitly consider (i) replay (retransmitting previously captured timing sequences), (ii) relay/wormhole (forwarding signals to emulate being in-region), (iii) anchor compromise (malicious transmitter emitting attacker-chosen timing), and (iv) collusion (multiple off-region receivers combining observations). Practical mitigations include time-bounded sessions (short validity windows tied to $T_{\text{net}}$ cadence), per-session nonces folded into the key derivation, and transmitter authentication of the timing sequence (e.g., cryptographic tags over the intended schedule). These steps align with adversarial-evaluation practices commonly used in robust security research: explicitly enumerating attacker capabilities, then stress-testing the system under those capabilities rather than relying on a single threat model.
How should $T_{\text{slot}}$ and the authorized region be selected in a principled way?
The work already links $T_{\text{slot}}$ to region size through timing tolerance (e.g., meters divided by $c$), and uses 256 slots corresponding to byte values. A principled selection should incorporate measurement noise (clock jitter effects) and application requirements (acceptable false-reject rate). One robust approach is to treat slot selection as an optimization: choose $T_{\text{slot}}$ to minimize off-region key success subject to an on-region decoding success constraint. Inspired by privacy-utility tuning in other security work (e.g., differential-privacy budget selection), $T_{\text{slot}}$ can be viewed as a “physical tolerance budget” that must be justified empirically.
What are the scalability limits (latency/throughput) of a timing-encoded key channel?
The design transmits 32 SHA bytes using a 33-packet schedule (one reference plus 32 differences). This fixed overhead means key distribution cost is non-trivial if the system refreshes keys frequently or serves many receivers. Two ways to improve scalability: (i) transmit fewer bits per session by deriving multiple content keys from a single in-region “seed” (key hierarchy), and (ii) add redundancy selectively (error-correcting encoding across bytes) only when channel conditions degrade. These improvements mirror ideas from resilient distributed systems and robustness research: amortize expensive secure setup, and add redundancy only where the failure modes concentrate.
@@ 7 Conclusion
By carefully scheduled UWB packet timing, this location-dependent cryptosystem enforces that decryption is possible only where the sender intends. Legitimate receivers within the authorized zone reconstruct the correct key and transparently decrypt audio content, while receivers outside this zone, even if they capture all packets, observe shifted timing patterns that yield incorrect keys. The system operates without ever exposing the password in human- or machine-readable form on the wire and initiates decryption without explicit user action, reducing both operational friction and key-handling risk.
@@
8 Acknowledgments
We gratefully acknowledge the support and sponsorship of Mr. Mike Ciholas and the guidance of our Design Advisor, Mr. Justin Bennett. We also thank our Academic Advisor, Dr. Donald Roberts, and Software Advisor, Mr. Tim DeBaillie, for their valuable technical feedback and support throughout the system design and implementation.